# : A Benchmark for Data-Efficiency in Long-Horizon Indoor Mobile Manipulation Robotics

### v6

@@ IV-A Baseline Models
To evaluate and establish baselines for future benchmarking, we test both learning-based and neuro-symbolic systems. Given the growing popularity of imitation learning, particularly BC, we benchmark two state-of-the-art (SOTA) BC models. This provides insight into the data efficiency of current BC algorithms.
@@
Additionally, we benchmark a neuro-symbolic system that integrates foundation models with task and motion planning (TAMP). The foundation model modules are trained on large-scale visual and textual data, but the system overall requires no robot demonstration training. Thus, we evaluated it zero-shot to assess whether non end-to-end methods can achieve greater data efficiency. By comparing its performance to learning-based approaches, this evaluation provides insight into whether such methods can achieve comparable or superior results, offering guidance for future development of data-efficient MoMa methodologies. Finally, we assess the performance of a random agent to establish a lower-bound baseline, wherein actions are uniformly sampled.
RT-1. Robotics Transformer (RT-1) [4] is a BC model designed for MoMa, leveraging a transformer architecture [31] to process natural language instructions and RGB image inputs into discretized actions. Its training on MoMa tasks, along with its status as an open-source, SOTA learning-based MoMa model, makes it particularly suitable for benchmarking on as a baseline.
@@
To adapt RT-1 for the ManipulaTHOR robot, we modified its action space to align with the robotâs capabilities. For task navigation, the robot base operates using discrete actions, while continuous deltas (discretized into 256 bins) are utilized for body rotation and manipulation.
MotionGlot-MoMa. MotionGlot (MG) [32] addresses the problem of varying action and observation dimensions in multi-embodiment models [33]. Leveraging training principles from language models, MG proposes a training procedure for motion generation across multiple embodiments as a next-token prediction problem.
@@
We adapt MG for MoMa by leveraging âs demonstrations consisting of paired RGB observations and actions, calling it MG-MoMa. MG-MoMa is trained to learn a language-conditioned model that maps those observations to corresponding actions. This process involves two stages: First, a VQ-VAEÂ [34] is employed to learn a discrete latent codebook for the images, forming the observation vocabulary, while discrete binning is applied to tokenize the actions. In the second stage, both the observation and action vocabularies are integrated with the text vocabulary from GPT-2Â [35]. This unified vocabulary enables action generation akin to text generation, utilizing the instruction-tuning template from MG-MoMa for text command-based action execution.
@@
LIMP. Language Instruction grounding for Motion Planning (LIMP) [36] is a neuro-symbolic system for pick-and-place MoMa integrating Linear Temporal Logic (LTL), Large Language Model (LLM), VLM, and novel TAMP algorithms to process natural language instructions into executable TAMP plans. By translating instructions into LTL formulas enriched with Composable Referent Descriptors [36], LIMP dynamically generates context-aware specification translations and compiles them into finite-state automata for verifiably correct execution. Unlike learning-based methods, LIMP requires no training on robot demonstrations, allowing it to be evaluated zero-shot. This enables a comparison of its data efficiency against BC models, providing insight into whether such systems are a promising direction for developing more data-efficient approaches to MoMa.
@@
To evaluate data efficiency, we assess how well models generalize to novel scenarios using Scene Generalization and Task Generalization, two widely used experimental setups [4, 15]. These evaluations are essential for benchmarking data efficiency because they test whether models can learn effectively from limited data and apply that knowledge to new scenarios. Scene Generalization reflects real-world deployment challenges where robots must perform tasks in previously unseen environments with different layouts and/or objects, while Task Generalization assesses adaptability to new tasks within seen settings. Together, these experiments quantify a modelâs data efficiency by isolating generalization challenges at both the scene and task levels, providing insights into how well models leverage limited training data to achieve robust performance.
@@
Data Split & Generalization. To implement these evaluations, we conduct experiments using carefully designed data splits on the simulation dataset, highlighted in Table II. For Scene Generalization, we employ a -fold cross-validation approach, where each of the five scenes in the dataset is held out as the test scene in turn. The held-out test set consists of all trajectories from the designated scene, while the training and validation sets are derived from the remaining four scenes. Specifically, of the trajectories from each of the four remaining scenes are included in the training set, and are allocated to the validation set, following a stratified split to preserve data distribution. During each fold, we train and hyperparameter tune a separate model specific to the held-out test scene. This process ensures that we assess the modelâs ability to generalize to entirely unseen environments while maintaining experimental controls. The rotation of held-out scenes allows for a comprehensive evaluation of Scene Generalization across the entire dataset.
@@
For Task Generalization, the data split is consistent across all five scenes. The training set comprises of the trajectories from each scene, the validation set includes , and the test set includes the remaining . Unlike Scene Generalization, the held-out test set in this experiment evaluates the modelâs ability to generalize to novel tasks rather than unseen environments. Hyperparameter tuning is performed, resulting in a single model optimized for Task Generalization on the test set.
@@
Training. To benchmark the data efficiency of modern BC methods, we train the RT-1 and MG-MoMa models from scratch without any pretrained parameters on the simulation tasks. This approach evaluates how well these models learn from our modest yet realistically sized dataset, providing a measure of their data efficiency specifically for long-horizon MoMa tasks. Both models are trained using the Scene Generalization and Task Generalization experiments, ensuring their capabilities are evaluated under varying conditions.
@@
Fine-Tuning. We fine-tune RT-1 and MG-MoMa with pretrained parameters and evaluate their performance on the simulation tasks. Empirical analysis showed that for RT-1, fine-tuning with some frozen model layers led to better performance than not freezing any, so this approach was used for the final experiment. In contrast, for MG-MoMa, the unfrozen model performed better, leading to the use of fully trainable parameters. Using this setup, we conducted the Task Generalization experiment to measure how much pretrained MoMa knowledge impacts the modelsâ ability to adapt efficiently to new MoMa tasks.
@@
Metric. To evaluate model performance, we use a task success rate metric that quantifies the completion of long-horizon MoMa tasks. Each task consists of four sequential subtasks: (1) navigating to the target object, (2) grasping the object, (3) transporting the object to the goal location, and (4) placing the object at the goal. Each successful subtask contributes one point, with a maximum of four points per task. The success rate is calculated by averaging the subtask scores within each task and then averaging across all tasks. This metric captures both navigation and manipulation performance, providing insight into a modelâs ability to execute multi-step long-horizon tasks.
@@ V-A Results
Scene Generalization. The success rates for the Scene Generalization experiment, detailed in Table III, are exceptionally low, with an overall average of just and across all scenes for RT-1 and MG-MoMa respectively. This demonstrates significant challenges in generalizing to novel scenes with unseen room layouts, furniture configurations, and object placements for our long-horizon MoMa tasks. Both models outperform the random agent, demonstrating at least a minimal capacity to learn something from the data.
Scenes 1 and 3 exhibited relatively higher success rates compared to the others for RT-1. This can be attributed to a relatively high proportion of evaluation tasks in these scenes involving objects that were positioned directly in front of or adjacent to the robotâs initial spawn location, reducing the complexity of navigation and manipulation required. In contrast, Scene 2 for RT-1 exhibited the lowest performance, likely due to the robot spawning in an enclosed and isolated room where many tasks necessitated navigating out of the room to retrieve the target object, significantly increasing task difficulty. Overall, MG-MoMa demonstrates performance that is near RT-1, which is noteworthy given that MG was not originally intended for MoMa tasks.
Task Generalization. Table IV presents the results of our Task Generalization experiments, which assess the modelâs ability to adapt to new tasks in a familiar environments. Notably, both RT-1 and MG-MoMa show slightly higher overall performance compared to the Scene Generalization experiments. This improvement is likely due to the larger training dataset size (see Table II) and the availability of all scenes during training. Moreover, while the performance of the baselines is equal for the seen settings, RT-1 outperforms MG-MoMa by on the unseen environments. Meanwhile, the random agent showed no difference in performance between seen and unseen tasks, as expected.
Task Generalization with Fine-Tuning. Interestingly, RT-1âs unseen task performance was marginally lower than the fully trained modelâs, as shown in Table IV. Similarly, MG-MoMa showed lower performance on seen tasks. This may mean that the pretrained parameters biased the models toward their original training tasks and away from effectively adapting to the unique challenges of , especially since the original models were trained in the real world.
@@
LIMP. The results presented thus far show that learning-based models (BC in particular) face significant challenges in addressing the complexities of âs long-horizon MoMa tasks. To explore an alternative paradigm, we benchmarked the LIMP system on the Task Generalization experiment.
The evaluation was conducted in real-world environments using Spot, assessing tasks, including a mix of tasks originally specified for both simulated and real-world environments. Since LIMP was originally designed for real-world execution, this mixed evaluation approach was necessary to also test its performance on simulation-specified tasks. For these simulation-designed tasks, real-world setups were constructed to approximate their simulated counterparts. This enabled us to evaluate LIMPâs versatility across both task types, providing insights into its adaptability to different environmental contexts. As shown in Table IV, LIMP achieved a relatively impressive average success rate of on unseen tasks, vastly outperforming RT-1 and MG-MoMa without any training on robot demonstrations, making it relatively more data efficient.
One key factor contributing to LIMPâs success is its ability to generate instruction-conditioned semantic maps of the environment. These maps provide the robot with precise information about the locations of objects and regions of interest, eliminating the need to search and enabling the interpretation of fine-grained semantics specified in language-conditioned tasks. Additionally, the maps allow the robot to utilize robust motion planning algorithms for navigation, which are less susceptible to covariate shift compared to the mapless action decoding used in BC models like RT-1. This robustness is critical for the tasksâ long-horizon navigation aspect, where the inherent large observation and state spaces exacerbate the challenges faced by learning-based methods.
Another significant advantage of LIMP is its incorporation of multimodal foundation models, which endow it with extensive world knowledge to reason about task requirements and perceptual inputs in a zero-shot manner. This capability enables it to handle the semantic and contextual complexities of our benchmark tasks without requiring task-specific training. While it demonstrates considerable promise, it still falls short of perfect execution, indicating room for improvement in handling the intricacies of our long-horizon MoMa tasks. Overall, LIMP represents a promising direction for future MoMa development, blending the strengths of learning-based approaches with those of classical planning.
@@ V-B Ablations
Impact of Dataset Size. ALOHA Unleashed [37], a bimanual manipulation system trained on K demonstrations, only achieved a success rate of for certain tasks, showing that even with large datasets, high performance is not guaranteed. Similarly, Lin et al. [38] found that increasing demonstrations generally improves manipulation performance, though with diminishing returns. Since our benchmark involves MoMaâwith larger action and observation spaces due to added navigationâand uses far fewer trajectories, substantial performance improvements likely require more data. To investigate this, we pose the following questions: Are the baseline models data inefficient? Will scaling the data improve their performance? To answer these, we investigate how performance scales when progressively larger subsets of the simulation dataset (, , , and ) are used for training in the Task Generalization experiment. This evaluates whether these models can effectively utilize limited data or require more, testing their data efficiency and scalability for long-horizon MoMa tasks.
@@
The ablation results, summarized in Figure 2, show a positive upward trend in performance for RT-1 (represented by the pink line) as the training data increases, albeit a weak one. This trend is evident in its average rate of change (RoC). This shows that scaling the data does improve RT-1âs overall performance. A similar trend is observed in the language, vision, and multimodal communities [39, 40, 41], making it encouraging to see this pattern emerge in the context of long-horizon MoMa tasks as well. Our results reflect that these BC models are data inefficient for long-horizon MoMa tasks, as the fewer the data, the worse it generally performs. This raises questions about the source of inefficiency, such as model architecture, explored in the following section.
@@
Impact of Architecture. To assess whether data inefficiency stems from the architecture, we evaluate and compare different architectures. While MG-MoMa was already benchmarked and provides some architectural comparison with RT-1, it is not sufficiently different since it also uses a transformer [31]. To explore distinct alternatives, we replace RT-1âs transformer with alternate architectures while keeping all other components, such as encoders, unchanged to ensure a controlled comparison. Specifically, we replace the transformer with an LSTM [42], which we refer to as RL-1. Likewise, we incorporate the Mamba architecture [43], leveraging its ability to handle long sequences [43], and call this RM-1. To understand how these alternative architectures perform with increasing data, we evaluate both RL-1 and RM-1 under the Task Generalization setup using the same subset-based evaluation as previously conducted. By comparing their trends/RoCs and full dataset results against RT-1âs, we aim to determine whether data inefficiency is tied to architectural limitations or not.
@@
As shown in Figure 2, both RL-1 and RM-1 exhibit upward trends, indicating that they did not learn more efficiently than RT-1 and that the data inefficiency is spanning across different architectures. Identifying the exact causes of data inefficiency requires further investigation, which is beyond the scope of this work and is left for the community to explore in the future.

### v7

@@ IV-A Baseline Models
To evaluate the benchmark and establish its initial baselines, we test four MoMa systems. Given the growing popularity of BC, we benchmark two state-of-the-art (SOTA) BC models. This provides insight into the data efficiency of current BC algorithms.
@@
Additionally, we benchmark a neuro-symbolic system that integrates foundation models with task and motion planning (TAMP). The foundation models are pretrained on large-scale visual and textual data, but the system overall requires no robot demonstration training. Thus, we evaluated it zero-shot to assess whether non-end-to-end methods can achieve greater data efficiency. By comparing its performance to learning-based approaches, this evaluation provides insight into whether such methods can achieve comparable or superior results, offering guidance for future development of data-efficient MoMa methodologies. Finally, we assess the performance of a random agent to establish a lower-bound baseline, wherein actions are uniformly sampled.
@@
Robotics Transformer. RT-1 [4] is a MoMa BC model. It processes natural language instructions and RGB image inputs into discretized actions by leveraging a transformer [31]. Its status as an open-source, SOTA learning-based MoMa model makes it suitable for benchmarking on as a baseline. To adapt RT-1 for the ManipulaTHOR robot, we modified its action head to align with the robotâs capabilities. For task navigation, the robot base operates using discrete actions, while continuous deltas (discretized into 256 bins) are utilized for body rotation and manipulation.
@@
MotionGlot-MoMa. MotionGlot (MG) [32] addresses the problem of varying action and observation dimensions in multi-embodiment models [33]. Leveraging training principles from language models, MG proposes a training procedure for motion generation across multiple embodiments as a next-token prediction problem. We adapt MG for MoMa by leveraging âs demonstrations consisting of paired RGB observations and actions, calling it MG-MoMa. MG-MoMa is trained to learn a language-conditioned model that maps those observations to corresponding actions. This process involves two stages. First, a VQ-VAE [34] learns a discrete latent codebook for the images, forming the observation vocabulary, while discrete binning is applied to tokenize the actions. Second, both the observation and action vocabularies are integrated with the text vocabulary from GPT-2 [35]. This unified vocabulary enables action generation akin to text generation, utilizing the instruction-tuning template from MG-MoMa for text command-based action execution.
@@
LIMP. Language Instruction grounding for Motion Planning (LIMP) [36] is a neuro-symbolic system for pick-and-place MoMa integrating Linear Temporal Logic (LTL), a Large Language Model (LLM), VLM, and novel TAMP algorithms to process natural language instructions into executable TAMP plans. By translating instructions into LTL formulas enriched with Composable Referent Descriptors [36], LIMP dynamically generates context-aware specification translations and compiles them into finite-state automata for verifiably correct execution. Unlike learning-based methods, LIMP requires no training on robot demonstrations, allowing it to be evaluated zero-shot. This enables a comparison of its data efficiency against BC models, providing insight into whether such systems are a promising direction for developing more data-efficient MoMa approaches.
@@
To evaluate data efficiency, we assess the models on scene generalization and task generalization, two widely used experiment types [4, 15]. These evaluations are essential for benchmarking data efficiency because they test whether models can learn effectively from limited data and apply that knowledge to novel scenarios. We define Scene Generalization as robots performing tasks in unseen environments with different layouts and/or objects, while Task Generalization as the ability to adapt to new tasks within seen settings. Together, these experiments quantify a modelâs data efficiency by isolating generalization challenges at both the scene and task levels, providing insights into how well models leverage limited training data to achieve robust performance.
@@
Data Split & Generalization. To implement these evaluations, we conduct experiments using carefully designed data splits on the simulation dataset, highlighted in Table II. For Scene Generalization, we employ a -fold cross-validation approach, where each of the five scenes in the dataset is held out as the test scene in turn. The held-out test set consists of all trajectories from the designated scene, while the training and validation sets are derived from the remaining four scenes. Specifically, of the trajectories from each of the four remaining scenes are included in the training set, and are allocated to the validation set, following a stratified split to preserve data distribution. During each fold, we train and hyperparameter-tune a separate model specific to the held-out test scene. This process ensures that we assess the modelâs ability to generalize to entirely unseen environments while maintaining experimental controls. The rotation of held-out scenes allows for a comprehensive evaluation of Scene Generalization across the entire dataset.
@@
For Task Generalization, the data split is consistent across all five scenes. The training set comprises of the trajectories from each scene, the validation set includes , and the test set includes the remaining . Unlike Scene Generalization, the held-out test set in this experiment is fixed and evaluates the modelâs ability to generalize to novel tasks rather than unseen environments. Hyperparameter tuning is performed, resulting in a single model optimized for Task Generalization on the test set.
@@
Training. We train RT-1 and MG-MoMa without pretrained parameters on the simulation tasks. This evaluates how well they learn from our realistically sized dataset, providing a measure of their data efficiency specifically for long-horizon MoMa tasks. Both models are trained using the Scene Generalization and Task Generalization experiments.
@@
Fine-Tuning. We fine-tune RT-1 and MG-MoMa with pretrained parameters on the simulation tasks. Our empirical analysis showed that for RT-1, fine-tuning with some frozen layers led to better performance than not freezing any, so this approach was used for the final experiment. In contrast, for MG-MoMa, the unfrozen model performed better, leading to the use of fully trainable parameters. This discrepancy may stem from MG-MoMa needing to override task-irrelevant biases from its original non-MoMa pretraining, necessitating updates across all layers. In contrast, since RT-1 was already pretrained on MoMa tasks, it may retain useful representations relevant to tasks in certain layers. Using this setup, we conducted the Task Generalization experiment to measure the extent to which pretrained MoMa knowledge impacts the modelsâ ability to adapt efficiently to new MoMa tasks.
@@
Metric. We use a task success rate metric that quantifies the completion of long-horizon MoMa tasks. Each task consists of four sequential dependent subtasks: (1) navigating to the target object, (2) grasping the object, (3) transporting the object to the goal location, and (4) placing the object at the goal. Each successful subtask contributes one point, with a maximum of four points per task. The success rate is calculated by averaging the subtask scores within each task and then averaging across all tasks. This captures both navigation and manipulation performance, providing insight into a modelâs ability to execute long-horizon tasks.
@@ V-A Results
Scene Generalization. The success rates for the Scene Generalization experiment, detailed in Table III, are exceptionally low, with an overall average of just and across all scenes for RT-1 and MG-MoMa, respectively. This demonstrates significant challenges in generalizing to novel scenes with unseen room layouts, furniture configurations, and object placements for our long-horizon MoMa tasks. Both models outperform the random agent, demonstrating at least a minimal capacity to learn something from the data.
Scenes 1 and 3 exhibited relatively higher success rates compared to the others for RT-1. This can be attributed to a relatively high proportion of evaluation tasks in these scenes involving objects that were positioned directly in front of or adjacent to the robotâs initial spawn location, reducing the complexity of navigation and manipulation required. In contrast, Scene 2 for RT-1 exhibited the lowest performance, likely due to the robot spawning in an enclosed and isolated room where many tasks necessitated navigating out of the room to retrieve the target object. Additionally, Scene 2 contained the highest number of objects compared to the other scenes (see Figure II). These factors combined significantly increased task complexity. Overall, MG-MoMa demonstrates performance that is near RT-1, which is noteworthy given that MG was not originally intended for MoMa tasks.
Task Generalization. Table IV presents the results of our Task Generalization experiments. Notably, both RT-1 and MG-MoMa show slightly higher overall performance compared to the Scene Generalization experiments. This improvement is likely due to the larger training dataset size (see Table II) and the availability of all scenes during training. Moreover, while the performance of the baselines is equal for the seen settings, RT-1 outperforms MG-MoMa by on the unseen environments. Meanwhile, the random agent showed no difference in performance between seen and unseen tasks, as expected.
Task Generalization with Fine-Tuning. RT-1âs unseen task performance was marginally lower than the fully trained modelâs, as shown in Table IV. Overall, seen and unseen results were similar, indicating that the models did not overfit. However, MG-MoMa exhibited relatively lower performance on seen tasks compared to unseen tasks when fine-tuned. This may mean that the pretrained parameters biased the models toward their original training tasks and away from effectively adapting to the unique challenges of , especially since the original models were trained on real-world data.
@@
LIMP. The results presented thus far show that learning-based models, BC in particular, face significant challenges in addressing the complexities of âs long-horizon MoMa tasks. To explore an alternative paradigm, we benchmarked the LIMP system on the Task Generalization experiment.
The evaluation was conducted in real-world environments using Spot, as LIMP was originally designed for real-world execution. We utilized tasks, drawn from a mix of simulated and real-world tasks. This mixed evaluation was necessary to test its performance on simulation-specified tasks as well. This enabled us to evaluate LIMPâs versatility across both task types, providing insights into its adaptability to different environmental contexts. For the simulation-designed tasks, real-world setups were constructed to approximate their simulated counterparts. As shown in Table IV, LIMP achieved a relatively impressive average success rate of on unseen tasks, vastly outperforming RT-1 and MG-MoMa without any training on robot demonstrations, making it relatively more data efficient.
One key factor contributing to LIMPâs success is its ability to generate instruction-conditioned semantic maps of the environment. These maps provide the robot with precise information about the locations of objects and regions of interest, eliminating the need to search and enabling the interpretation of fine-grained semantics specified in language-conditioned tasks. Additionally, the maps allow the robot to utilize robust motion planning algorithms for navigation, which are less susceptible to covariate shift compared to the mapless action decoding used in BC models like RT-1. This robustness is critical for the tasksâ long-horizon navigation aspect, where the inherent large observation and state spaces exacerbate the challenges faced by learning-based methods. Another significant advantage of LIMP is its incorporation of multimodal foundation models, which endow it with world knowledge to reason about task requirements and perceptual inputs in a zero-shot manner. This capability enables it to handle the semantic and contextual complexities of our benchmark tasks without requiring task-specific training.
While LIMP demonstrates considerable promise, it still falls short of perfect execution, indicating room for improvement in handling the intricacies of our long-horizon MoMa tasks. In particular, revealed that LIMP exhibits an object-centric bias; it lacks an explicit understanding of spatial constructs such as rooms (e.g., it can navigate to a âcupâ but not to a âkitchenâ). It also lacks a notion of the starting location, rendering it unable to interpret instructions that require returning to its point of origin. Moreover, due to its reliance on a VLM, its performance is constrained by the VLMâs perceptual limitations, which in this case led to failures in recognizing or locating target objects. Overall, LIMP represents a promising direction for future MoMa development, blending the strengths of learning-based approaches with those of classical planning.
@@ V-B Ablations
Impact of Dataset Size. ALOHA Unleashed [37], a bimanual manipulation system trained on K demonstrations, only achieved a success rate of for certain tasks, showing that even with large datasets, high performance is not guaranteed. Similarly, Lin et al. [38] found that increasing demonstrations generally improves manipulation performance, though with diminishing returns. Since our benchmark involves MoMaâwith larger action, state, and observation spaces due to added navigationâand uses far fewer trajectories, substantial performance improvements likely require more data. To investigate this, we pose the following questions: Are the baseline models data inefficient? Will scaling the data improve their performance? To answer these, we investigate how performance scales when progressively larger subsets of the simulation dataset (, , , and ) are used for training in the Task Generalization experiment. This evaluates whether these models can effectively utilize limited data or require more, testing their data efficiency and scalability for long-horizon MoMa tasks.
@@
The ablation results, summarized in Figure 2, show a positive upward trend in performance for RT-1 (represented by the pink line) as the training data increases, albeit a weak one. This trend is evident in its average rate of change (RoC). This shows that scaling the data does improve RT-1âs overall performance. This pattern is consistent with what is observed in the language, vision, and multimodal communities [39, 40, 41]. Our results reflect that these BC models are data inefficient for long-horizon MoMa tasks, as the less data, the worse they generally perform. This raises questions about the source of inefficiency, such as model architecture.
@@
Impact of Architecture. To assess whether data inefficiency stems from the architecture, we evaluate and compare different architectures. While MG-MoMa was already benchmarked and provides some architectural comparison with RT-1, it is not sufficiently different since it also uses a transformer [31]. To explore distinct alternatives, we replace RT-1âs transformer with alternate architectures while keeping all other components, such as encoders, unchanged to ensure a controlled comparison. Specifically, we replace the transformer with an LSTM [42], and refer to the model as RL-1. Likewise, we incorporate the relatively new Mamba architecture [43], and call this RM-1. To understand how these alternative architectures perform with increasing data, we evaluate both RL-1 and RM-1 under the Task Generalization setup using the same subset-based evaluation previously conducted. By comparing their trends/RoCs, subsets, and full dataset results against RT-1âs, we aim to determine whether data inefficiency is tied to architectural limitations.
@@
As presented in Figure 2, both RT-1 and RL-1 exhibit upward trends with similar RoCs. This indicates that RL-1 did not learn more efficiently than RT-1. However, RM-1 maintained a relatively constant trend and outperformed RT-1 across all subsets and the full dataset. This finding shows that architecture could be a contributing factor in the observed data inefficiency associated with -type tasks. Identifying more concrete and exact causes of data inefficiency requires further investigation, which is beyond the scope of this work and is left for the community to explore.